OpenStudy (anonymous):

find the equation of the tangent to the curve y=X^(x^2) - X^(Pi) where x=1 find the equation of the tangent to the curve y=X^(x^2) - X^(Pi) where x=1 @Mathematics

OpenStudy (anonymous):

Take log of y and then differentiate... (dy/dx) at x=1 will give the slope of the tangent...

OpenStudy (anonymous):

when you say dy/dx at x=1 do you mean to plug in 1 for all x after i get the derivative of the equation?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

is the slop -10.6063

OpenStudy (anonymous):

you can also find y at x=1... Thus you know the slope and one point in the line which will help you to find the equation...

OpenStudy (anonymous):

what was your dy/dx

OpenStudy (anonymous):

is it not (1-pi)?

OpenStudy (anonymous):

yes yes thats the slope right? then how do i find the y to plug in for point slope form

OpenStudy (anonymous):

you have the equation of the curve... put x=1 in it... that'll give you the point required...

OpenStudy (anonymous):

that is y=X^(x^2) - X^(Pi) at x=1

OpenStudy (anonymous):

= 0

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so you've got a point and the slope... go on...

OpenStudy (anonymous):

can i leave it at y-0=1-pi(x-1)

OpenStudy (anonymous):

y-0=(1-pi)(x-1) don't forget the brackets...

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

you may write it in the form y=mx+c if you like...

OpenStudy (anonymous):

y=-piX +X+pi-1

OpenStudy (anonymous):

y=(1-pi)x-(1-pi)

OpenStudy (anonymous):

thank you soo much!

OpenStudy (anonymous):

you're welcome...

OpenStudy (anonymous):

A piece of advice: you should've calculated the derivative by hand....

OpenStudy (anonymous):

yea thats the problem i dont know how to do that one

OpenStudy (anonymous):

do you want me to explain?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

In y=X^(x^2) - X^(Pi) the problem is with the first term... I believe that you know how to differentiate x^(pi)...

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

let y=p-q where p=X^(x^2) and q=X^(Pi)

OpenStudy (anonymous):

taking log of p log(p)=x^2*log(x) then differentiate (1/p)dp/dx=x^2/x+2xlog(x) dp/dx=p[x+2xlog(x)] dp/dx=x^(x^2)*[x+2xlog(x)]

OpenStudy (anonymous):

y=dp/dx-dq/dx

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

(1/p)dp/dx=x^2/x+2xlog(x)

OpenStudy (anonymous):

so derivative of log(x) = 2xlog(x)

OpenStudy (anonymous):

nope... derivative of log(x) is 1/x

OpenStudy (anonymous):

p is the product of x^2 and log(x) so you've to use the product rule.

OpenStudy (anonymous):

\[d(x^2*\log(x))/dx=x^2*d(\log(x))/dx + \log(x)*d(x^2)/dx\]

OpenStudy (anonymous):

where is the 1/p from? (1/p)dp/dx=x^2/x+2xlog(x)

OpenStudy (anonymous):

where did the x^2 go? dp/dx=p[x+2xlog(x)]

OpenStudy (anonymous):

x^2/x=x